# Simple Yet Effective: Extracting Private Data Across Clients in Federated Fine-Tuning of Large Language Models

### v1

@@ Appendix A Acronyms List
  • GDPR - General Data Protection Regulation [European Union, 2016]
  • CCPA - California Consumer Privacy Act [State of California, US, 2018]
  • CPIS - Chinese Information Security Technology: Personal Information Security Specification (GB/T 35273-2020) [Standardization Administration of China (2020), SAC]
  • Singapore PDPC - Personal Data Protection Commission (Singapore) [Personal Data Protection Commission, Singapore, 2012]
@@ B.1 Data Extraction Attack
Early research on training data extraction attacks has broadly categorized them into untargeted and targeted attacks [Research, 2022, Yu et al., 2023]. Untargeted extraction aims to recover any memorized training samples without specifying a target [Lukas et al., 2023], whereas targeted extraction attempts to reconstruct specific training samples, often by providing a known prefix and recovering the remaining content [Carlini et al., 2021]. The latter type, often referred to as Verbatim Data Extraction, has become a standard approach for evaluating memorization in LLMs [Carlini et al., 2023, Dong et al., 2024] and for detecting potential data contamination [Dong et al., 2024]. We briefly outline the core methodology of verbatim data extraction below.
@@
In practice, model outputs may not exactly replicate the original suffix but can still be lexically close. To accommodate this, a similarity-based metric such as Edit Distance [Levenshtein, 1965] is often employed. A sample is deemed extracted if the similarity score between $g_{i}$ and $b_{i}$ exceeds a predefined threshold $t$. By computing this similarity-based extraction score across all samples in a dataset $D$, one can quantify the model’s memorization behavior or assess its vulnerability to training data extraction attacks.
@@ B.2 Federated Learning
Federated Learning (FL) is a solution to address data isolation issues [Yang et al., 2019], where a central server and multiple clients collaborate to complete the training process. A key feature of FL is that the training datasets are stored locally on each client and remain invisible to other clients. FL is commonly used in industrial scenarios where each client represents an independent organization, such as hospitals collaborating to train a medical model without combining their datasets due to legal restrictions or business competition. Federated Learning enables the training of stronger models compared to training on data from a single client alone.
@@
To solve this problem, many federated optimization algorithms have been proposed, such as FedAVG [McMahan et al., 2023] and FedProx [Li et al., 2020]. Typically, these algorithms consist of two alternating phases: local updating and central aggregation. In the local updating phase, each client independently optimizes the global model using its own dataset. In the central aggregation phase, the server aggregates the models from the clients using an aggregation algorithm, obtaining a global model, which is then sent back to each client for the next round of local updating. A typical procedure of federated learning is illustrated in Algorithm 1.
Appendix C A Framework of Federated Learning Algorithms
@@ Appendix E Data Preprocessing
Previous works [Zhang et al., 2023, Yue et al., 2023] have used these datasets for LLM and FedLLM research. In this work, we use the processed datasets from these prior studies and further curate the data for our experiments. We applied the following preprocessing steps to prepare the datasets:
@@
Appendix F Details of Dataset
@@
Appendix G Details about PII Machine Annotation
Table 6: Categorization of Personally Identifiable Information (PII) Types in Our Labeling Standards
@@ Appendix G Details about PII Machine Annotation
Appendix H Implementation Details
H.1 Federated Dataset Partitioning.
We use the preprocessed and labeled datasets (see Section 3) for our experiments, splitting the data into training and testing sets. In the federated learning setup, we simulate a system with five clients. The testing set remains global, while the training set is heterogeneously partitioned across the clients using a balanced Non-IID distribution (see Acronyms List A). To achieve this, We employ a clustering-based method [Li et al., 2023] for partitioning, where a language encoder first generates embeddings, which are then clustered using K-means. Finally, a Dirac distribution with $\alpha=0.5$ is applied to create a label-skewed partitioning [Guo et al., 2024], ensuring each client receives a comparable number of samples.
H.2 Hardware.
All experiments are conducted on a single Nvidia A6000 GPU with 48GB of memory, with a float precision of bf16.
H.3 Experiment Procedure
H.3.1 Federated Utility Fine-Tuning
We begin by performing federated fine-tuning of the LLM[Zhang et al., 2023, Wu et al., 2025] on the partitioned dataset, adapting it to the legal tasks. The fine-tuning is conducted using the OpenFedLLM framework Ye et al. [2024b]. We set the total number of FL rounds to 10 and use FedAVG as the aggregation algorithm.
Each client performs multi-task fine-tuning by mixing all local tasks and applying a unified prompt template, as illustrated in Figure 6, following the approach in Raffel et al. [2023]. In each round of federated learning, the client fine-tunes the received global model for one epoch using parameter-efficient fine-tuning (PEFT) techniques of LoRA [Hu et al., 2021]. The learning rate is set to 3e-4 with a linear decay schedule. The maximum input sequence length is 3072 tokens. We use a batch size of 1 and apply gradient accumulation with a factor of 8. The LoRA configuration is set to $r=16$ and $\alpha=32$.
@@ H.3.2 PII Extraction
For Frequency-Prioritized Prefix Sampling, we construct $\text{Set}(\text{SUP}(P_{c}))$ from the aforementioned $P_{c}$, and sort it in descending order of prefix frequency (as described in Section 4.2.1). The model $\theta$ is then queried using prefixes in this frequency-descending order. Although we do not explicitly define a frequency threshold $\sigma_{a}$, we sweep the prefix budget $B$ exponentially in base 10. Because $\text{Set}(\text{SUP}(P_{c}))$ is sorted by decreasing frequency, this sweep over $B$ implicitly corresponds to sweeping $\sigma_{a}$ from +$\infty$+ ∞ to 1.
H.3.3 Latent Association Fine-tuning
@@
Appendix I Resutls of FedLLM Utility Fine-tuning
<### Input>
{{Task Input}}
<### Output>
{{Task Output}}
Figure 6: Unified Utility Fine-tuning Template for All Tasks.
Table 8: Utility Performance over Different Tasks.
FedAvg Qwen1-8B 50.0 14.2 50.0 37.5 90.0
FedAvg Baichuan2-7B 57.6 42.4 50.0 33.3 89.5
Non-FL Qwen1-8B 50.0 18.9 50.0 40.8 87.0
Appendix J Ablation Study Results
Table 9: Coverage rate (CR) of attacks for different attacker–victim client pairs with a prefix budget $B=10000$. Prefixes are randomly sampled from the attacker’s corresponding set $P_{c}$. "–" denotes cases where the attacker and victim are the same client (not applicable).
Attacker ID Victim ID Average
0 1 2 3 4
0 - 10.91% 12.89% 10.93% 11.88% 11.66%
1 11.97% - 12.41% 11.46% 12.35% 12.05%
2 12.56% 11.39% - 11.65% 12.74% 12.09%
3 12.07% 10.82% 12.04% - 11.99% 11.73%
4 12.26% 11.36% 13.25% 11.21% - 12.02%

### v2

@@ 6 Conclusion
Limitations
This work investigates the privacy risks of FedLLMs using a legal-domain dataset. Future research can extend our proposed methods to other sensitive domains such as healthcare and finance, where privacy concerns are equally critical. Additionally, there is a need for further exploration of defense mechanisms that can preserve the privacy of FedLLMs while maintaining their performance.
Ethics Statement
This paper presents PII extraction attacks on federated fine-tuned LLMs to expose potential privacy risks. While designed for research and defense purposes, such methods could be misused to recover sensitive user data in real-world FL systems. We conduct all experiments on legal datasets with anonymized PII, and highlight the need for stronger safeguards in FedLLM deployments.
Acknowledgements
We thank Guanzhong Chen and Yukun Zhang for their help with dataset annotation. All annotators were properly compensated for their contributions.
@@ Appendix A Acronyms List
  • GDPR - General Data Protection Regulation (European Union, 2016)
  • CCPA - California Consumer Privacy Act (State of California, US, 2018)
  • CPIS - Chinese Information Security Technology: Personal Information Security Specification (GB/T 35273-2020) (Standardization Administration of China (2020), SAC)
  • Singapore PDPC - Personal Data Protection Commission (Singapore) (Personal Data Protection Commission, Singapore, 2012)
@@ B.1 Data Extraction Attack
Early research on training data extraction attacks has broadly categorized them into untargeted and targeted attacks (Research, 2022; Yu et al., 2023). Untargeted extraction aims to recover any memorized training samples without specifying a target (Lukas et al., 2023), whereas targeted extraction attempts to reconstruct specific training samples, often by providing a known prefix and recovering the remaining content (Carlini et al., 2021). The latter type, often referred to as Verbatim Data Extraction, has become a standard approach for evaluating memorization in LLMs (Carlini et al., 2023; Dong et al., 2024) and for detecting potential data contamination (Dong et al., 2024). We briefly outline the core methodology of verbatim data extraction below.
@@
In practice, model outputs may not exactly replicate the original suffix but can still be lexically close. To accommodate this, a similarity-based metric such as Edit Distance (Levenshtein, 1965) is often employed. A sample is deemed extracted if the similarity score between $g_{i}$ and $b_{i}$ exceeds a predefined threshold $t$. By computing this similarity-based extraction score across all samples in a dataset $D$, one can quantify the model’s memorization behavior or assess its vulnerability to training data extraction attacks.
@@ B.2 Federated Learning
Federated Learning (FL) is a solution to address data isolation issues (Yang et al., 2019), where a central server and multiple clients collaborate to complete the training process. A key feature of FL is that the training datasets are stored locally on each client and remain invisible to other clients. FL is commonly used in industrial scenarios where each client represents an independent organization, such as hospitals collaborating to train a medical model without combining their datasets due to legal restrictions or business competition. Federated Learning enables the training of stronger models compared to training on data from a single client alone.
@@
To solve this problem, many federated optimization algorithms have been proposed, such as FedAVG (McMahan et al., 2023) and FedProx (Li et al., 2020). Typically, these algorithms consist of two alternating phases: local updating and central aggregation. In the local updating phase, each client independently optimizes the global model using its own dataset. In the central aggregation phase, the server aggregates the models from the clients using an aggregation algorithm, obtaining a global model, which is then sent back to each client for the next round of local updating. A typical procedure of federated learning is illustrated in Algorithm 1.
B.3 Federated Large Language Models (FedLLMs)
In current research at the intersection of federated learning and large language models, the term FedLLM predominantly refers to federated fine-tuning of pre-trained LLMs (Bai et al., 2024; Wang et al., 2024b; Ye et al., 2024a), rather than federated pre-training. This focus arises from both practical and technical considerations.
Federated pre-training is rarely necessary, as large-scale pre-training typically relies on publicly available general-purpose corpora (such as web text or encyclopedias) that do not contain sensitive information and therefore do not require federated sharing. Moreover, federated pre-training would entail transmitting extremely large model parameters across institutions, incurring prohibitive communication costs and conflicting with high-efficiency training practices like data parallelism and optimized operators, making deployment infeasible in practice.
By contrast, federated fine-tuning is essential in privacy-sensitive domains. Many applications of LLMs in vertical fields—such as judicial documents, electronic health records, and banking customer data—rely on restricted information that cannot be centralized due to legal or institutional constraints (e.g., the Data Security Law or the Personal Information Protection Law). In these settings, federated fine-tuning allows each institution to adapt a shared pre-trained model locally, achieving a balance between model performance and data privacy. This approach has already demonstrated tangible value across multiple domains.
For example, in the judicial domain (Zhang et al., 2023), courts can fine-tune a common LLM on local case repositories without sharing sensitive records, enabling cross-court tasks such as statute matching and case similarity analysis. In healthcare (Ali et al., 2025), hospitals can locally fine-tune models on specialty data, producing comprehensive medical LLMs capable of supporting structured record processing and disease risk prediction. And in finance (Sha, 2024), banks can fine-tune models on transaction data to detect fraud and assess credit risk without violating privacy regulations.
B.4 Preliminary Assessment of Verbatim Data Extraction Risks in FedLLMs
To examine the memorization behavior of FedLLMs and evaluate their potential risks of leaking sensitive information, we conduct a preliminary experiment simulating a verbatim data extraction attack. The results are referenced in the main paper (Section 1) to empirically motivate our study.
We adapt the experimental setup from (Dong et al., 2024) to the federated setting, where an attacker is assumed to possess prefix fragments of the training data from all participating clients and attempts to recover the subsequent suffix tokens. For each training sample, we extract a prefix from the original sequence and query the trained model to generate a continuation. The generated suffix is compared against the ground truth using Edit Distance (ED) (Levenshtein, 1965), where a lower ED indicates stronger memorization. Specifically:
  • ED = 0 indicates the model has perfectly memorized and reproduced the suffix;
  • ED values are capped at 50, as we restrict suffixes to a maximum of 50 tokens.
We perform the attack on the global models aggregated after 10 rounds of federated training. To ensure a comprehensive assessment, we consider three popular FL algorithms—FedAvg (McMahan et al., 2023), FedProx (Li et al., 2020), and Scaffold (Karimireddy et al., 2020)—each under both IID and Non-IID data distributions. Two baseline settings are also included:
  • Centralized: All client data is pooled and the model is fine-tuned in a conventional non-federated manner;
  • Untrained: The base model is evaluated without any fine-tuning.
Figure 7 summarizes the results across five downstream tasks. Our key observations are:
Figure 7: Edit Distance results of verbatim data extraction attacks after 10 training rounds. We evaluate six federated configurations (FedAvg/FedProx/Scaffold $\times$ IID/Non-IID) and report the mean and standard deviation. Lower values indicate stronger memorization. Centralized and untrained models serve as baselines.
  • FedLLMs consistently exhibit higher ED scores (i.e., lower memorization) than centralized models, suggesting that the FL aggregation process reduces susceptibility to verbatim extraction.
  • However, compared to untrained models, FedLLMs still show non-negligible memorization, with noticeably lower ED scores, indicating partial leakage of training data.
These findings highlight a trade-off between collaborative model training and privacy preservation, and they serve as the motivation for our in-depth investigation of privacy risks in FedLLMs.
Appendix C Federated Learning Framework
Algorithm 1 outlines a general framework for Federated Learning (FL), where a central server coordinates multiple clients to collaboratively train a global model without sharing local data. At each round, the server distributes the current model to all clients, each of which performs local updates based on its private data and sends the updated parameters back. The server then aggregates the received updates to produce a new global model.
@@ Appendix E Data Preprocessing
Previous works (Zhang et al., 2023; Yue et al., 2023) have used these datasets for LLM and FedLLM research. In this work, we use the processed datasets from these prior studies and further curate the data for our experiments. We applied the following preprocessing steps to prepare the datasets:
@@
Appendix F Supplementary Dataset Statistics and Analysis
Table 7 summarizes the basic statistics of the five datasets used in our experiments. Each dataset corresponds to a different downstream task for fine-tuning the model.
Figure 8 presents the document frequency distribution of the 2017 VxPII instances extracted from the model trained on the masked dataset (see Section 5.4). Most VxPII exhibit low frequency, indicating that PII masking significantly reduces memorization.
@@ Appendix F Supplementary Dataset Statistics and Analysis
Appendix G Prompt Template and Utility Fine-tuning Results for FedLLMs
Figure 9 shows the unified prompt template used for all federated utility fine-tuning tasks. Table 8 reports the evaluation results across multiple tasks, comparing different federated learning algorithms and base models.
<### Input>
{{Task Input}}
<### Output>
{{Task Output}}
Figure 9: Unified Utility Fine-tuning Template for All Tasks.
Table 8: Utility Performance over Different Tasks.
FedAvg Qwen1-8B 50.0 14.2 50.0 37.5 90.0
FedAvg Baichuan2-7B 57.6 42.4 50.0 33.3 89.5
Non-FL Qwen1-8B 50.0 18.9 50.0 40.8 87.0
Appendix H Additional PII Label Distribution Results
Figure 10 illustrates the label distribution of FedLLM-exclusive victim PII extracted by Qwen1-8B. This result corresponds to the experiment described in Section 5.4.
Figure 11 presents the label distribution of deduplicated victim-exclusive PII instances extracted by the Baichuan2-7B model.
Figure 11: Label distribution of deduplicated victim-exclusive PII instances extracted by the Baichuan2-7B model (without LAFt, using prefix set $P_{c}$). This figure complements Figure 3 in the main text, which presents the corresponding results for Qwen1-8B.
Appendix I Machine Annotation Standards for PII Labeling
This section provides details on the machine annotation protocol we use to identify PII in our dataset. Table 9 defines our categorization schema, which includes seven major categories and their corresponding fine-grained subtypes. To ensure annotation consistency and scalability, we utilize a templated prompting approach for automated PII labeling. Figure 12 shows the machine annotation prompt used to instruct the LLM annotator. The prompt dynamically incorporates category definitions and format constraints to standardize the output.
Table 9: Categorization of PII Types in Our Labeling Standards
@@ Appendix I Machine Annotation Standards for PII Labeling
Appendix J Details of Human Evaluation for PII Annotation
To validate the quality of the machine-generated PII annotations, we recruited four Chinese-speaking students with foundational knowledge of Chinese law to manually annotate PII on a selected subset of the dataset. Prior to annotation, all annotators underwent thorough training on the annotation guidelines and usage of the Label Studio tool. The instructions provided to annotators are detailed in Figure 13, while Figure 14 illustrates the annotation interface used. All annotators were fairly compensated upon completion of their tasks.
The human evaluation results, reported in terms of precision, recall, and F1 score, are summarized in Table 10, indicating high agreement both in exact span matching and in combined span-and-label matching, confirming the reliability of the machine annotations.
# **PII Annotation Guidelines for Labelers**
## **1. Task Objective**
**Core Task**: Proofread legal texts to accurately identify and annotate **Personally Identifiable Information (PII)**. Each annotation task includes:
1. **Localization**: Mark the exact character offsets of each PII instance in the text;
2. **Categorization**: Assign each PII instance to the appropriate **major category (7 total)** and **minor category (36 total)**, ensuring precise classification.
## **2. PII Category System**
| Major Category | Minor Categories |
| Personal Basic Information | Name, Birthday, Address, Gender, Ethnicity, Family Relationship, Age, Nationality, Personal Phone Number |
...(omitted)...
## **3. Annotation Workflow and Standards**
### **Step-by-Step Process**
1. **Read the Full Text**: Understand the context to detect all potential PII entities;
2. **Sentence-by-Sentence Annotation**: For each PII instance, annotate its **start position**, **text span**, and corresponding **major + minor category**;
3. **Special Cases**: For ambiguous expressions (e.g., "a certain district of a certain city"), determine PII status based on contextual clues.
### **Annotation Guidelines**
* **Accuracy**: Ensure all annotated content is verifiably present in the text. Avoid false positives or over-labeling;
* **Support Channel**: If any uncertain cases arise during annotation, promptly reach out to the *Annotation Support Team* for clarification.
Figure 13: Markdown-style guideline for PII annotation, covering task objectives, taxonomy, and labeling procedures.
Figure 14: Human annotation interface in the Label Studio tool for PII labeling. Annotators are familiar with the Label Studio environment and are instructed to label PII spans based on predefined PII categories. Machine-generated labels are provided as references to assist the human annotators.
@@
Appendix K Experiment Implementation Details
K.1 Federated Dataset Partitioning.
We use the preprocessed and labeled datasets (see Section 3) for our experiments, splitting the data into training and testing sets. In the federated learning setup, we simulate a system with five clients. The testing set remains global, while the training set is heterogeneously partitioned across the clients using a balanced Non-IID distribution (see Acronyms List A). To achieve this, We employ a clustering-based method (Li et al., 2023) for partitioning, where a language encoder first generates embeddings, which are then clustered using K-means. Finally, a Dirac distribution with $\alpha=0.5$ is applied to create a label-skewed partitioning (Guo et al., 2024), ensuring each client receives a comparable number of samples.
K.2 Hardware and Computation Budget
All experiments are conducted on a single NVIDIA A6000 GPU with 48 GB of memory, using bfloat16 precision. Most sampling-based attack experiments are completed within 200 GPU hours.
K.3 Experiment Procedure
K.3.1 Federated Utility Fine-Tuning
We begin by performing federated fine-tuning of the LLM (Zhang et al., 2023; Wu et al., 2025) on the partitioned dataset, adapting it to the legal tasks. The fine-tuning is conducted using the OpenFedLLM framework (Ye et al., 2024b). We set the total number of FL rounds to 10 and use FedAVG (McMahan et al., 2023) as the aggregation algorithm.
Each client performs multi-task fine-tuning by mixing all local tasks and applying a unified prompt template, as illustrated in Figure 9, following the approach of Raffel et al. (2023). In each round of federated learning, the client fine-tunes the received global model for one epoch using parameter-efficient fine-tuning (PEFT) techniques of LoRA (Hu et al., 2021). The learning rate is set to 3e-4 with a linear decay schedule. The maximum input sequence length is 3072 tokens. We use a batch size of 1 and apply gradient accumulation with a factor of 8. The LoRA configuration is set to $r=16$ and $\alpha=32$.
@@
For Frequency-Prioritized Prefix Sampling, we construct $\text{Set}(\text{SUP}(P_{c}))$ from the aforementioned $P_{c}$, and sort it in descending order of prefix frequency (as described in Section 4.2.1). The model $\theta$ is then queried using prefixes in this frequency-descending order. Although we do not explicitly define a frequency threshold $\sigma_{a}$, we sweep the prefix budget $B$ exponentially in base 10. Because $\text{Set}(\text{SUP}(P_{c}))$ is sorted by decreasing frequency, this sweep over $B$ implicitly corresponds to sweeping $\sigma_{a}$ from $\infty$ to 1.
K.3.3 Latent Association Fine-tuning